### v2

# COMI-LINGUA: Expert Annotated Large-Scale Dataset for Multitask NLP in Hindi-English Code-Mixing

Abstract
We introduce COMI-LINGUA, the largest manually annotated Hindi-English code-mixed dataset, comprising 125K+ high-quality instances across five core NLP tasks: Matrix Language Identification, Token-level Language Identification, POS Tagging, Named Entity Recognition (NER), and Machine Translation. Each instance is annotated by three bilingual annotators, yielding over 376K expert annotations with strong inter-annotator agreement (Fleiss’ Kappa 0.81). The rigorously preprocessed and filtered dataset covers both Devanagari and Roman scripts and spans diverse domains, ensuring real-world linguistic coverage. Evaluation reveals that closed-source LLMs significantly outperform traditional tools and open-source models. Notably, one-shot prompting consistently boosts performance across tasks, especially in structure-sensitive predictions like POS and NER, highlighting the effectiveness of prompt-based adaptation in code-mixed, low-resource settings. COMI-LINGUA is publicly available at this URL111 https://anonymous.4open.science/r/CodeMixing/.
COMI-LINGUA: Expert Annotated Large-Scale Dataset for Multitask NLP in Hindi-English Code-Mixing
Rajvee Sheth, Himanshu Beniwal, Mayank Singh LINGO, Indian Institute of Technology Gandhinagar, India Correspondence: lingo@iitgn.ac.in
1 Introduction
Code-mixing is the blending of multiple languages within a single utterance—a pervasive phenomenon in multilingual societies, especially on social media platforms (Jamatia et al., 2020; Srivastava and Singh, 2020). Over half of the world’s population is bilingual or multilingual and frequently uses mixed-language expressions in digital communication (Grosjean, 2021). In the Indian context, Hindi-English (Hinglish) code-mixed text is particularly widespread and presents significant computational challenges due to orthographic complexity, frequent language switches, and script variation between Devanagari and Roman forms (Bali et al., 2014; Takawane et al., 2023; Thara and Poornachandran, 2018). A characteristic example is:
Kal mujhe hai, but
will be an issue,
where Hindi and English tokens co-occur and certain English words like “office” and “traffic” may appear in Devanagari script.
(English Translation: “Tomorrow I have to go to the office, but traffic will be an issue.”)
Despite growing interest, current Hinglish datasets have critical limitations: (1) a predominant focus on Roman script, ignoring natural script variation (Begum et al., 2016; Bali et al., 2014; Srivastava et al., 2020), (2) limited scale and coverage (Srivastava and Singh, 2021a; Kumar et al., 2018; Tiwari et al., 2024; Kartik et al., 2024), (3) insufficient task diversity within single datasets (Aguilar et al., 2020; Khanujaa et al., 2020; Bohra et al., 2018; Khanuja et al., 2020), and (4) reliance on synthetic data generation rather than human annotation (Chatterjee et al., 2022; Srivastava and Singh, 2021c; Kartik et al., 2024; Sravani and Mamidi, 2023).
To address these limitations, we present a novel comprehensive dataset COMI-LINGUA (COde-MIxing and LINGuistic Insights on Natural Hinglish Usage and Annotation) that advances Hinglish code-mixing research. The key contributions include:
Curation of the largest publicly available Hinglish dataset (376K manually annotated instances), released under a CC-BY-4.0 license, capturing real-world code-mixing behavior across both Roman and Devanagari scripts. Each instance is annotated by three annotators across five key NLP tasks: matrix language identification, token-level language identification, part-of-speech tagging, named entity recognition, and machine translation.
Robust benchmarking of state-of-the-art multilingual LLMs (mLLMs), including both open-weight and closed-source models—alongside traditional NLP tools, under two inference paradigms: zero-shot and one-shot in-context learning.
In-depth error analysis of mLLMs on code-mixed tasks, uncovering critical limitations such as misclassification of English borrowings in Devanagari script, context truncation, overfitting in one-shot settings, prompt mimicry, repetitive or hallucinated outputs, and practical deployment barriers like API usage constraints—highlighting persistent challenges in script-aware and context-sensitive language modeling.
2 Related Work
Code-mixing—the blending of multiple languages in a single utterance—poses major challenges for NLP due to its structural variability (Srivastava and Singh, 2021a). This is especially true for Hinglish, given their distinct scripts and syntax (Bali et al., 2014). Progress is hindered by the lack of large, annotated datasets, as collecting and labeling such data remains costly and labor-intensive (Srivastava and Singh, 2021a).
Language Identification is a foundational task in code-mixed NLP. Multiple approaches have been developed to detect language boundaries within mixed-language sequences, including statistical models, CRFs, and deep learning-based techniques (Shekhar et al., 2020; Singh et al., 2018a; Gundapu and Mamidi, 2018; Molina et al., 2016). These efforts have paved the way for improved preprocessing and downstream modeling of code-mixed data.
Part-of-Speech Tagging A variety of annotated datasets have been introduced for POS tagging in code-mixed contexts. Singh et al. (2018b) and Vyas et al. (2014) developed corpora from Twitter and Facebook, respectively, while Pratapa et al. (2018) generated synthetic datasets for evaluating bilingual word embeddings. Sequiera et al. (2015) experimented with various machine learning algorithms, and Chatterjee et al. (2022) introduced PACMAN, a large-scale synthetic POS-tagged dataset that achieved state-of-the-art performance in code-mixed POS tagging tasks.
Named Entity Recognition in code-mixed text has seen significant progress through both resource development and model improvements. Dowlagar and Mamidi (2022) showed that leveraging multilingual data enhances NER accuracy, while Ansari et al. (2019) created cross-script datasets using Wikipedia. Transformer-based approaches and meta-embeddings have also been effective in improving NER for Indian code-mixed data (Priyadharshini et al., 2020).
Machine Translation for code-mixed content remains a growing research area. Dhar et al. (2018) and Srivastava and Singh (2020) developed parallel corpora for Hinglish code-mixed sentences, while Hegde and Lakshmaiah (2022) proposed translation models using transliteration and pseudo-translation, achieving competitive results in the MixMT shared task at WMT 2022.
Benchmarking and Evaluation Frameworks Several benchmark datasets have been introduced to evaluate NLP systems on code-mixed tasks. LinCE (Aguilar et al., 2020) provides a comprehensive benchmark covering 11 corpora and 4 language pairs. GLUECoS (Khanuja et al., 2020) demonstrated the benefits of fine-tuning multilingual models on code-switched datasets across multiple tasks. Emotion and sentiment annotation efforts, such as the Hinglish Twitter corpus by Vijay et al. (2018), the L3Cube-HingCorpus (Nayak and Joshi, 2022), and the emotion-annotated SentiMix dataset by Ghosh et al. (2023) further support affective computing in code-mixed settings.
Despite ongoing efforts, standardized benchmarks for evaluating LLMs on diverse Hinglish code-mixed tasks—such as acceptability judgments, syntactic fluency, and translation fidelity—remain limited. Existing benchmarks are often narrow in scope and rely on synthetic or small-scale data. To address this, we curate the largest high-quality, human-annotated dataset for training and evaluating LLMs on a broad range of Hinglish code-mixed phenomena. It serves as both an evaluation suite and a diagnostic tool to advance multilingual and code-mixed language understanding research.
3 The COMI-LINGUA dataset
3.1 Raw Dataset Curation
We curated raw data from publicly accessible and licensed platforms spanning diverse categories such as news, politics, entertainment, social events, sports, and informational content, with a focus on the Indian subcontinent. Sources included prominent news portals and official digital archives, detailed in Appendix 17. The collected content was cleaned using regex-based preprocessing to remove noise such as advertisements, HTML tags, and footers, and then segmented into individual sentences. A Code-Mixing Index (CMI, Das and Gambäck (2014)) was computed for each sentence, and only those sentences with a CMI score 9 were retained to ensure a substantial degree of code-mixing. Given the under-representation of mixed Devanagari-Roman script samples in existing datasets, we also collected supplementary data to enhance coverage and linguistic diversity. This includes enriching the dataset by incorporating additional Hinglish code-mixed samples from prior works (Srivastava and Singh, 2020; Gupta et al., 2023; Singh et al., 2018c) and from Hugging-Face222https://huggingface.co/datasets/pardeep/youtube-vidoes-transcripts-hindi-english/.
3.2 Dataset Processing
The preprocessing pipeline was designed to enhance the quality and neutrality of the corpus through rigorous noise reduction techniques. To ensure the dataset was both clean and relevant, we removed duplicate instances, hate speech, and abusive content. Sentences containing offensive or inappropriate language were identified and filtered out using established profanity and hate speech detection tools, including thisandagain333https://github.com/thisandagain/washyourmouthoutwithsoap/blob/develop/README.md and Hate-Speech-Detection-in-Hindi444https://github.com/victorknox/Hate-Speech-Detection-in-Hindi/blob/main/README.md.
At the token level, additional preprocessing steps were applied. Sentences with fewer than five tokens were discarded to eliminate non-informative content such as fragments, abbreviations, emojis, and filler phrases—commonly arising from typing errors or social media discourse. Examples of such removed content include: #GuessTheSong’, during dinner’, and ‘@enlightenedme bas ek hi’. Further data refinement was conducted during the manual annotation process (see Section 3.4 for more details).
3.3 Data Annotation
To annotate the Hinglish code-mixed corpus, we employed Commentator (Sheth et al., 2024), a robust annotation framework specifically designed for multilingual code-mixed text.
The annotation was carried out by a team of three graduate-level experts proficient in both Hindi and English. All annotators possess prior experience with social media content and demonstrate strong programming capabilities, along with familiarity with using version control systems. These competencies contributed to a systematic, efficient, and reproducible annotation process. The annotators were recruited specifically for this project and were compensated at a rate of approximately $1.64 per hour. The funding for the annotation work was provided through a government-sponsored initiative; the compensation adheres to standard remuneration practices considered appropriate for the annotators’ qualifications and the economic context of their country of residence.
We selected five diverse annotation tasks, balancing well-established tasks with high reliability and underexplored challenges. Annotators followed detailed guidelines with examples to ensure consistency and clarity across tasks (Appendix A.3, Figure 1). The tasks are:
1.
Token-level Language Identification (LID): In this task, each token in the dataset was assigned one of three possible language labels: English (en), Hindi (hi), or Other (ot). Initial language tags were generated using Microsoft’s Language Identification Tool555https://github.com/microsoft/LID-tool, which served as a baseline for further manual refinement. As shown in Figure 1, each token is assigned a language tag.
2.
Matrix Language Identification (MLI): Each sentence is annotated with a Matrix Language, which identifies the dominant language governing the grammatical structure of the sentence. In code-mixed text, even when multiple languages are interspersed, one language typically dictates the syntactic and morphosyntactic framework of the utterance. Figure 1 showcases a sentence annotated with its matrix language.
3.
Named Entity Recognition (NER): In the NER task, each token in a sentence is annotated with a label from a predefined set of entity types outlined in Table 1. These include conventional categories, such as Person, Location, Organization, Date/Time, and GPE (Geo-Political Entities), as well as social media-specific types like Hashtags, Mentions, and emojis. An instance of annotated entities across different types is shown in Figure 1. This allows the annotation schema to comprehensively capture the diversity and informality often observed in code-mixed social media text.
Entity Type Description Person Names of individuals Location Non-political physical locations Organization Institutions or companies Date/Time Temporal expressions (e.g., dates) GPE Geo-Political Entities Hashtags Words prefixed by ‘#’ Mentions User mentions prefixed by ‘@’ Emoji Emoticons conveying emotions Table 1: Named entity types and their descriptions in our annotation schema. -
4.
Part-of-Speech (POS) Tagging: Each token in the code-mixed dataset was annotated with a Part-of-Speech (POS) tag selected from the Universal POS tagset proposed by Singh et al. (2018b). The tagset, summarized in Table 2, was chosen for its language-agnostic design, enabling consistent annotation of Hindi and English words in a single sentence—an essential feature for handling code-mixed content effectively. A representative example is presented in Figure 1. Initial predictions for POS tags were generated using the CodeSwitch NLP library666https://github.com/sagorbrur/codeswitch, which supports multilingual code-mixed data and provides pre-trained models suitable for tagging noisy, informal text commonly found on social media platforms.
POS Tag Description NOUN Common nouns PROPN Proper nouns VERB Verbs in all tenses and moods ADJ Adjectives describing nouns ADV Adverbs modifying verbs ADP Adpositions PRON Pronouns DET Determiners CONJ Conjunctions PART Grammatical particles PRON_WH Wh-pronouns PART_NEG Negative particles NUM Numerals and cardinal numbers X Typos, abbreviations, punctuation Table 2: Part-of-speech tags and their descriptions used in our annotation schema. -
5.
Machine Translation (MT): This task involves constructing parallel translations for code-mixed sentences into three distinct formats: (i) Romanized Hindi, (ii) Devanagari-script Hindi, and (iii) standard English. The goal is to facilitate a multilingual Hinglish sentence to align with its respective translations across scripts and languages. A representative translation instance across the three formats is shown in Figure 1. Initial translation predictions were generated using the LLaMA 3.3 language model777https://github.com/meta-llama/llama-models/blob/main/models/llama3_3/MODEL_CARD.md.
For all tasks, we used state-of-the-art NLP tools or LLMs for automated pre-annotation, generating initial labels based on task-specific criteria. Expert annotators then refined these outputs through manual post-editing. This two-stage process ensured high-quality annotations while improving consistency and speeding up dataset creation.
3.4 Manual Data Refinement
During the annotation phase, the dataset underwent iterative refinement to ensure quality and consistency, guided by annotator feedback on instances to be excluded (see Table 6 in Appendix A.2). Sentences were removed if they (i) were monolingual English or Hindi, (ii) lacked relevant linguistic tags or named entities, contained no meaningful content, or merged multiple instances into one, or (iii) included languages other than Hindi and English, which were beyond the scope of this study. This refinement process was crucial for preserving corpus integrity and ensuring that the final dataset consisted solely of high-quality Hinglish code-mixed text. The Raw and Filtered columns in Table 3 represent the number of original instances provided for initial annotation and the final number of instances retained after annotation, respectively. The difference between these values corresponds to instances flagged by annotators as not satisfying the manual annotation criteria.
3.5 Annotation Efforts and Quality
The manual annotation process involved substantial human effort across all tasks, particularly in refining the outputs of automated tools. For example, for the LID task, each annotator reviewed 504,102 tokens and flipped an average of 95,670 tokens—approximately 19% of the original predictions. In the POS task, 63,002 of 427,941 tokens were corrected, indicating a 15% flip rate. Similarly, for the NER task, each annotator modified about 98,760 out of 538,160 tokens, translating to 18% manual corrections. For the MLI task, no initial predictions were provided, leading to 100% of the sentences being annotated. To assess annotation reliability, we computed inter-annotator agreement (IAA) using Fleiss’ Kappa (Fleiss, 1971), a standard metric for evaluating consistency among multiple annotators on categorical labels (Hallgren, 2012). All classification tasks achieved Fleiss’ Kappa scores above 0.817, indicating substantial to near-perfect agreement (Table 3). As machine translation is a generative task, IAA was not calculated. While not a direct measure of quality, the final dataset retains a high level of code-mixing, with an average CMI exceeding 14 across tasks, ensuring strong code-mixing.
Task | Raw | Filtered | IAA | CMI |
LID | 29,950 | 25,772 | 0.834 | 20.87 |
MLI | 29,950 | 25,772 | 0.976 | 20.87 |
POS | 27,229 | 24,598 | 0.817 | 21.60 |
NER | 26,929 | 24,913 | 0.852 | 14.38 |
MT | 26,727 | 24,558 | - | 17.07 |
Total / Avg. | 140,785 | 125,613 | 0.863 | 18.96 |
Model/Library | LID | POS | NER | MLI | MT | ||||||||||
claude-3.5-sonnet | 92.8 | 92.4 | 92.1 | 75.3 | 64.8 | 69.0 | 59.1 | 55.1 | 56.7 | 98.8 | 83.5 | 90.0 | 48.0 | 48.6 | 56.0 |
gpt-4o | 92.8 | 92.8 | 92.7 | 76.1 | 66.0 | 70.1 | 60.5 | 60.1 | 60.1 | 98.4 | 97.9 | 98.1 | 28.8 | 27.4 | 31.9 |
gemini-1.5-Flash | 82.9 | 40.4 | 47.9 | 73.4 | 62.4 | 66.5 | 44.2 | 44.2 | 43.8 | 98.8 | 21.4 | 33.7 | 48.1 | 28.9 | 56.9 |
LLaMA-3.3-instruct | 73.4 | 73.7 | 73.3 | 74.3 | 65.5 | 68.9 | 67.5 | 67.3 | 66.8 | 98.8 | 59.0 | 73.1 | 55.4 | 50.4 | 59.8 |
mistral-instruct | 54.5 | 39.0 | 42.4 | 10.2 | 6.72 | 7.78 | 65.1 | 41.5 | 50.2 | 98.1 | 58.7 | 72.3 | 23.5 | 5.36 | 18.05 |
command-a-03-2025 | 92.0 | 92.0 | 91.8 | 73.5 | 65.4 | 68.6 | 65.9 | 67.8 | 66.6 | 98.5 | 98.0 | 98.3 | 38.6 | 35.1 | 48.8 |
codeswitch | - | - | - | 89.1 | 87.8 | 88.2 | 81.6 | 83.1 | 81.2 | - | - | - | - | - | - |
Microsoft LID | 80.2 | 76.5 | 74.4 | - | - | - | - | - | - | - | - | - | - | - | - |
Model/Library | LID | POS | NER | MLI | MT | ||||||||||
claude-3.5-sonnet | 93.0 | 92.7 | 92.5 | 81.4 | 79.2 | 79.3 | 85.9 | 85.2 | 85.0 | 98.8 | 98.9 | 98.8 | 50.9 | 52.1 | 63.4 |
gpt-4o | 93.9 | 94.0 | 93.8 | 81.6 | 78.0 | 78.9 | 77.4 | 75.8 | 76.0 | 98.7 | 97.7 | 98.1 | 50.1 | 50.2 | 58.26 |
gemini-1.5-Flash | 80.2 | 76.5 | 74.4 | 72.9 | 64.6 | 68.0 | 66.5 | 67.5 | 66.0 | 98.4 | 40.4 | 56.4 | 44.4 | 42.9 | 66.1 |
LLaMA-3.3-instruct | 90.3 | 89.6 | 89.3 | 85.1 | 84.0 | 84.1 | 79.0 | 79.1 | 78.4 | 98.8 | 97.8 | 98.2 | 62.2 | 54.4 | 60.6 |
mistral-instruct | 72.1 | 70.0 | 70.1 | 77.3 | 66.9 | 69.8 | 65.5 | 44.4 | 52.6 | 98.3 | 88.1 | 92.7 | 30.0 | 19.5 | 18.5 |
command-a-03-2025 | 92.1 | 91.7 | 91.3 | 74.5 | 65.7 | 69.5 | 76.7 | 78.9 | 77.3 | 98.9 | 98.7 | 98.3 | 52.9 | 42.1 | 56.0 |
The COMI-LINGUA consists of 125,613 high-quality instances spanning five tasks, each independently annotated by three expert annotators, yielding a total of 376,839 annotations (see Table 7) . To our knowledge, it is the largest manually annotated code-mixed dataset to date. For each task, we provide two random splits: a test set of 5,000 instances and a training set comprising the remainder.
4 Experiments
4.1 Baseline Tools and LLMs
We conduct a comprehensive evaluation of existing tools and language models on the COMI-LINGUA Benchmark. Our experimental setup spans traditional NLP toolkits, state-of-the-art open-source LLMs, and proprietary commercial models. These systems are evaluated on their performance across five diverse Hinglish code-mixed NLP tasks, detailed in Section 3.3.
The traditional tools evaluated in this study include the Microsoft LID888https://github.com/microsoft/LID-tool designed for token-level language identification in multilingual text, and the codeswitch toolkit999https://github.com/sagorbrur/codeswitch, which provides a rule-based pipeline for annotating syntactic and semantic information in code-switched corpora. The open-source LLMs considered in our evaluation include mistral-instruct (7B) (Jiang et al., 2023) and llama-3.3-instruct (70B) (Touvron et al., 2023). In addition, we assess three commercial state-of-the-art systems: GPT-4o (Achiam et al., 2023), command-a-03-2025 (111B) (Cohere et al., 2025), Claude-3.5-Sonnet Anthropic (2024) and Gemini-1.5-Flash (Anil et al., 2023).
We create specific prompt templates for each task to generate accurate, task-aligned responses from LLMs. The prompt template includes a high-level description of the task, specific annotation or tagging rules, and illustrative examples where applicable. For each of our five tasks, we developed two prompt variants: a zero-shot version providing only task instructions and a one-shot version that includes a single demonstrative example with instructions. The prompts are presented as a system-level instruction, followed by the user-supplied test input (i.e., a code-mixed sentence or token sequence). The complete prompt template used for one task is detailed in Appendix B.
4.2 Evaluation Metrics
We employ a suite of standard evaluation metrics, appropriately chosen for each task’s nature. For token-level classification tasks—LID, POS, and NER—we report Precision (P), Recall (R), and the F1-score, computed at the macro level. For the MLI task, which is a sentence-level classification problem, we adopt the same classification metrics—P, R, and F1—computed on a per-sentence basis. For MT, we use the BLEU score (Papineni et al., 2002) to evaluate the quality of translated outputs. Given the multilingual nature of our dataset, BLEU is computed separately for each output format: for English, for Romanized Hindi, and for Devanagari Hindi. This disaggregated evaluation helps assess script-specific translation quality and is especially relevant given the transliteration variability in informal code-mixed text.
4.3 Evaluation Configurations
We evaluate model performance under two distinct inference paradigms: zero-shot and one-shot (specifically, 1-shot) in-context learning. Traditional NLP tools and libraries are inherently limited to zero-shot settings, as they rely on fixed rule-based or statistical models without the capability for contextual adaptation. In contrast, LLMs are evaluated under both zero-shot and 1-shot configurations to investigate their ability to generalize from instructions alone and to leverage minimal contextual supervision, respectively.
In the zero-shot setting, the prompt includes only task-specific instructions and formatting constraints without any illustrative examples. For the 1-shot setting, we augment the prompt with a single representative example demonstrating the input-output structure of the task. This example is carefully selected to reflect typical task behavior and is kept fixed across all evaluations to maintain consistency. Detailed illustrations of both prompt configurations are provided in Appendix B.
5 Results and Observations
Tables 4 and 5 present the empirical results obtained under the two experimental configurations: zero-shot and one-shot in-context learning, respectively. It is important to note that traditional tools such as codeswitch and Microsoft LID are limited in their task coverage; consequently, results for tasks not supported by these tools are omitted from the tables.
Traditional Tools vs. LLMs: The comparative analysis of traditional NLP tools and LLMs reveals clear distinctions in performance across code-mixed tasks. As shown in Table 4, traditional tools such as codeswitch and Microsoft LID demonstrate strong performance on specific tasks they were designed for, particularly POS and LID, respectively. For instance, codeswitch achieves the highest POS F1-score of 88.2, outperforming all LLMs in this task, while Microsoft LID attains a reasonable F1-score of 74.4 for LID. However, these tools exhibit significant limitations in task coverage; they do not support MLI, MT, or tasks involving complex reasoning or generation.
Open vs. Closed LLMs
The performance gap between proprietary (closed) and open-source LLMs is evident across both zero-shot and few-shot settings. In zero-shot mode, closed models such as gpt-4o and claude-3.5-sonnet dominate with top-tier results in most tasks. For example, gpt-4o achieves 92.7 F1 on LID and 98.1 F1 on MLI, while claude-3.5-sonnet reaches 92.1 F1 on LID and 90.0 F1 on MLI. However, when moving to a one-shot setting, open-source models like LLaMA-3.3-instruct start closing the gap. Its performance improves significantly: LID F1 rises from 73.3 to 89.3, POS tagging reaches 84.1 (even surpassing gpt-4o), and NER climbs to 78.4. MT performance also peaks at 62.2 for English, the highest across all models.
Zero vs. One-shot Inference
The transition from zero-shot to one-shot inference leads to notable performance improvements across most models and tasks. This is especially evident in complex tasks such as NER and MT, where providing one task-specific instance helps models disambiguate entities and manage code-mixed structures more effectively. For example, Claude-3.5-sonnet’s NER F1 score increases significantly from 56.7 in the zero-shot setting to 85.0 in the one-shot setting, while its for Devanagari Hindi translation improves from 31.9 to 63.4. gpt-4o similarly benefits, with NER performance rising from 60.1 to 76.0 and Devanagari improving from 31.9 to 58.26. Open models like LLaMA-3.3-instruct also see considerable gains, such as POS tagging jumping from 68.9 to 84.1 and English MT reaching 62.2. These results demonstrate that even minimal supervision through a single example can significantly enhance model performance on linguistically complex, low-resource, or code-mixed tasks. At the same time, tasks like MLI exhibit relatively modest gains, suggesting that more deterministic tasks benefit less from one-shot prompting. Overall, one-shot inference provides a practical and effective method to unlock the latent capabilities of LLMs in multilingual and code-mixed scenarios.
6 Challenges with Current LLMs
A consistent challenge across all models is the inability to accurately handle English borrowings written in Devanagari script—words like “” and “
” were frequently misclassified as Hindi, reflecting a gap in script-aware language identification. Another prominent issue is sentence truncation; longer code-mixed inputs often lead to incomplete or abruptly cut-off outputs, indicating that many models struggle to preserve context over extended sequences. In addition, models such as gemini-1.5-flash and mistral-instruct displayed repetitive generation patterns, producing redundant phrases within the same response. These models also occasionally injected subjective explanations into their outputs, despite clear instructions to extract objective information—for instance, adding interpretive statements when identifying the matrix language. A further concern is that several models tended to mirror patterns from the prompt rather than perform actual analysis, indicating shallow understanding and a tendency to copy input structures. Sentences with high grammatical or script variability posed yet another barrier, where many models, especially gemini-1.5-flash and mistral-instruct, failed to generate any output at all. Overfitting to examples also emerged as a concern, particularly in one-shot settings; models like gpt-4o and command-a-03-2025 occasionally produced outputs that mimicked example structures rather than responding appropriately to the test input. This over-reliance was particularly evident in tasks such as MLI and LID, where one-shot performance slightly declined. Additionally, some models exhibited hallucination behaviors, introducing entities not present in the input, which likely stems from overgeneralization in the presence of minimal supervision (See Table 8 in the Appendix).
7 Conclusion and Future Directions
LLMs often struggle with tasks like POS tagging, NER, and MT in Hinglish code-mixed text due to limited exposure to Indian multilingual data. This results in errors such as entity mislabeling and content hallucination, especially with structurally complex and script-variable inputs. The COMI-LINGUA dataset addresses these challenges by providing high-quality, richly annotated, and task-diverse code-mixed text. Future work includes expanding the dataset and adding tasks such as normalization, sentiment analysis, Q&A, summarization, and dialogue understanding, creating a stronger foundation for training and evaluating LLMs on complex scenarios.
Limitations
While this study offers valuable insights into the annotation and processing of Hinglish code-mixed text, several limitations warrant consideration:
1.
Language Pair Specificity: The findings derived from Hinglish code-mixed data may not generalize to other language pairs (e.g., Spanish-English), given differences in syntactic structure, sociolinguistic norms, and code-switching behavior.
2.
Demographic Bias: The use of a relatively small and homogeneous group of annotators may introduce demographic bias, potentially limiting the broader applicability and reliability of the acceptability ratings.
3.
Resource Constraints: Scaling this work to other code-mixed language pairs remains challenging due to the scarcity of high-quality annotated corpora and the limited availability of models capable of robustly handling diverse code-mixing phenomena.
Ethics Statement
We adhere to established ethical guidelines in the creation of our benchmark dataset and in the evaluation of existing LLMs for Hinglish code-mixed text. Data curation was carried out responsibly, with careful attention to the annotator’s well-being, informed consent, and workload management. We ensured that no personally identifiable information (PII) was included in the dataset, thereby maintaining user privacy and confidentiality. To mitigate potential biases, annotation protocols were designed to capture diverse linguistic phenomena and were reviewed iteratively. Our study promotes fairness and inclusivity in multilingual NLP by focusing on underrepresented code-mixed language scenarios. All datasets and models employed in this research are either publicly available or used in accordance with their respective licenses, such as Creative Commons.
Appendix A Appendix
A.1 Dataset Sources
For dataset collection, we implemented an article-wise scraping process that extracted high-quality data from diverse sources. News sources included NDTV101010https://ndtv.in/, ABP News111111https://www.abplive.com/, Zee News121212https://zeenews.india.com/hindi, News18131313https://hindi.news18.com/, TV9141414https://www.tv9.com/, and Aaj Tak.151515https://www.aajtak.in/ Digital platforms like X (formerly “Twitter”)161616https://x.com/ and YouTube171717https://www.youtube.com/ provided real-time discussions. Political channels from INC, BJP, and AAP were included, along with official sources such as Mann Ki Baat181818https://www.narendramodi.in/mann-ki-baat and Press Information Bureau (PIB)191919https://pib.gov.in/.
A.2 Examples of Noisy Text Instances in the Scrapped Code-Mixed Data
Table 6 Presents examples of challenging text patterns identified during manual annotation, including incomplete variants, ambiguous scripts, cross-article concatenation, and mixed-script forms. These were carefully reviewed and, in some cases, removed as part of our annotation methodology and quality assurance process to improve dataset consistency.
A.3 Annotation Guidelines for All Tasks
Each instance was annotated independently by all annotators without influence from model predictions or other annotator’s decisions.
Annotators were instructed to rely on contextual understanding to disambiguate code-mixed tokens, resolve ambiguity, and accurately assign labels.
Only the content explicitly present in the sentence was to be annotated; annotators were advised to avoid adding any inferred or assumed information.
Instances containing noise (e.g., incomplete fragments, junk tokens, or malformed words) were marked and excluded during preprocessing as per filtering heuristics as per Table6.
Consistent labeling was promoted using uniform tags and task-specific instructions during annotation training.
Annotators were encouraged to flag uncertain, ambiguous, or low-quality samples for further review.
Annotation disagreements were addressed using majority voting. In cases where no majority existed, a manual adjudication process was conducted to finalize the labels.
Appendix B Experimental Setup
B.1 Zero-Shot Prompt
Few-Shot Prompt
Category | Example Text |
Incomplete variant |
), floppy disk, hard disk drive, magnetic stripe card, relational database, SQL ![]() ![]() |
Ambiguous script |
Menu<br/>![]() ![]() |
Cross-article concatenation |
Mixed-script variant | @Strawberigloz he barobar naahi aahe, aaplich manasa aaplyala paathi sodtat. Aaplya itithasacha garva asla pahije. |
Task | Data Source (Hi-En) | Dataset Size | Script | QA | Annotators/Models |
LID | Facebook (Bali et al., 2014) | 1,062 | R & D | Yes | 3 |
Twitter Singh et al. (2018a) | 2,079 | R | Yes | 3 | |
Twitter (Swami et al., 2018) | 5,250 | R | Yes | Not mentioned | |
Twitter (Mave et al., 2018) | 5,567 | R | Yes | 3 | |
Facebook, Twitter, WhatsApp (Veena et al., 2018) | 3,071 | R | No | Embedding Model | |
Twitter (Joshi and Joshi, 2022) | 18,461 | R | No | Not mentioned | |
Twitter, YouTube, Press Releases, News (Ours) | 25,772 | R & D | Yes | 3 | |
POS | Twitter, Facebook (Sequiera et al., 2015) | 628 | R & D | No | 1 |
Facebook (Bali et al., 2014) | 1,062 | R & D | Yes | 3 | |
Twitter, Facebook (Jamatia et al., 2015) | 1,106 | R | No | 2 | |
Twitter (Singh et al., 2018b) | 1,190 | R | Yes | 3 | |
Synthetically generated (Chatterjee et al., 2022) | 51,118 | R & D | No | 0 | |
Existing Benchmarks (Kodali et al., 2022) | 55,474 | R | No | Trained POS tagger | |
Twitter, YouTube, Press Releases, News (Ours) | 24,598 | R & D | Yes | 3 | |
NER | Facebook (Bali et al., 2014) | 1,062 | R & D | Yes | 3 |
Twitter (Singh et al., 2018a) | 2,079 | R | Yes | 3 | |
Twitter (Bhargava et al., 2016) | 2,700 | R | No | Supervised algorithm | |
Twitter (Singh et al., 2018c) | 3,638 | R | Yes | 2 | |
Tourism, News (Murthy et al., 2022) | 108,608 | R & D | No | 1 | |
Twitter, YouTube, Press Releases, News (Ours) | 24,913 | R & D | Yes | 3 | |
MT | TED Talks, News, Wikipedia Kartik et al. (2024) | 2,787 | R & D | Yes | 2 |
Twitter, Facebook Srivastava and Singh (2021b) | 3,952 | R & D | Yes | 5 | |
Social Media Dhar et al. (2018) | 6,096 | R | Yes | 4 | |
Twitter, Facebook Srivastava and Singh (2020) | 13,738 | R | Yes | 54 (400 instances) | |
Existing Benchmarks Kunchukuttan et al. (2017) | 14,95,854 | R & D | No | PBSMT, NMT | |
Twitter, YouTube, Press Releases, News (Ours) | 24,558 | R & D | Yes | 2 | |
MLI | Twitter, Facebook (Sequiera et al., 2015) | 628 | R & D | No | 1 |
Facebook (Bali et al., 2014) | 1,062 | R & D | Yes | 3 | |
Social Media (Dhar et al., 2018) | 6,096 | R | Yes | 4 | |
Twitter, YouTube, Press Releases, News (Ours) | 25,772 | R & D | Yes | 3 |
Response Flaw Type | Example Behavior or Observation |
Script and entity Misidentification |
Words such as ![]() |
Sentence Truncation | Long-form code-mixed inputs lead to abrupt endings or incomplete generations (e.g., output stops mid-sentence despite ample context). |
In “Yes, we belong to this place – ![]() ![]() ![]() |
Repetitive Generation | Models like gemini-1.5-flash and mistral-instruct frequently exhibit repetitive generation patterns. For instance, they may produce outputs such as: “The second tagging is more accurate as it identifies ‘this’ as a determiner and ‘last’ as a quantity. repeating similar explanations or sentence fragments within the same response. |
Subjective Additions | Instead of remaining factual, models add speculative commentary (e.g., “en: The given text is in English. The hashtag "#MadeByGoogle" is also in English. ‘E’ (English).”). |
Prompt Mimicry | gpt-4o and command-r-plus mirror example formats from the prompt, failing to adapt to new inputs and instead mimicking example structure. Based on the given text, it is written in the Hindi language. Therefore, the matrix language label for this sentence is ‘h’. |
High-variance Failure | Inputs with abrupt transitions, broken grammar, or inconsistent scripts result in empty, irrelevant, or default responses. |
Example 1: lakhanOo: dr. apj abdul kalam bhArat kA 11veN rAShTrapati thE karoDhON bhAratiyON ke lIyE prEraNAdhA kA strOt thE, dR. apj abdul kalam | |
Example 2: This text does not contain any GPE, DATE, TIME, HASHTAG, EMOJI, or MENTION entities. | |
Hallucination | Models like mistral-instruct fabricate non-existent locations or attributes (e.g., inventing MATCH, VERSION, COUNTRY, PRODUCT, QUANTITY, or BUILDING categories not present in the input). |
Example 1: Note: The context is assumed to be empty in this example. If context information is available, it should be provided to improve the accuracy of entity tagging. | |
Example 2: ḍipṭī sīṭī mānejar lēmūēla rāṇḍolpha nē kahā, ’splāś pāḍ kā rōjānā sāf-sāfāī kamī huī hai |
Appendix C Computation Requirement and Budget
The experiments were conducted using API-based access to state-of-the-art Large Language Models (LLMs), including gpt-4o, Command R+ (command-a-03-2025) by Cohere, and claude-3.5-sonnet. The estimated monthly costs for API usage were approximately $200 for Claude-3.5-sonnet, $150 for Cohere, and $50 for gpt-4o, resulting in a total estimated cost of $400 per month.